### v1

@@ Abstract
This year’s centenary celebrations for the founding of quantum mechanics by Werner Heisenberg in 1925 also coincides with the thirtieth anniversary of the death of Rudolph Peierls in 1995. Peierls was Heisenberg’s student, perhaps his most illustrious one, well known to be the father of the atomic bomb, see Frank Close FrankClose1 . In writing this commentary on Peierls’s paper of Physics Today in 1991 Peierls1991 I thought it might be useful to recall some memories and discussions I had with Rudy, as we called him in Oxford, pertaining in particular to fundamental questions in quantum mechanics, as a tribute to both men and as a prelude to my views on his response to Bell Bell1990 . Heisenberg was of course famous for his discovery of matrix mechanics in 1925 and as a student Peierls had learnt a lot from him and his way of doing physics and also from Pauli who had many anecdotes which Peierls recalled to me on occasions over the dinner table. Concerning quantum mechanics, I was always interested in discussions with him on the Landau and Peierls (1930) Peierls-Landau1930 paper on the uncertainty in the measurement of fundamental quantities imposed by relativity (finite speed of light $c$), which is the Landau-Peierls expression for the relativistic Heisenberg uncertainty principle: $\Delta p\Delta t\sim\hbar/c$. This would imply amongst other things that a precise measurement for the momentum say of the electromagnetic field i.e. the photon (which by default is a relativistic object) would require an infinite amount of time. This result came as such a shock to Niels Bohr that both young men, Peierls and Landau, were detained in Copenhagen in 1930, unable to leave until Bohr was satisfied that a resolution has been found. In fact, it was not until 1950 that Bohr and Rosenfeld finally published their paper Bohr-Rosenfeld 1950 , which was a masterpiece on measurement analysis, which Bohr insisted must be conducted with classical apparatus. The psychological effects on both Peierls and Landau must have been deep and also on Heisenberg who later became involved Bohr-Rosenfeld 1950 . Heisenberg later changed his own earlier views on quantum mechanics and adopted the S matrix theory for particle interaction processes Heisenberg1938 , as the only description that has observable physical meaning. As for Peierls, my feeling was that he was completely broken down by Bohr (a very persistent and persuasive man) while Landau may not have been so. Landau never accepted quantum field theory as a logically consistent theory due to infinities but partly due to the Landau-Peierls results of 1930 Peierls-Landau1930 ; Landau-Lifshitz4 . Prior to this they had also together penned their paper on quantum electrodynamics in configuration space, proposing for the first time a Schrödinger equation for light quanta Peierls-Landau21930 . The paper inspired Enrico Fermi to use their methodology for the theory of beta decay.

### v2

@@ Abstract
This year’s centenary celebrations for the founding of quantum mechanics by Werner Heisenberg in 1925 also coincides with the thirtieth anniversary of the death of Rudolph Peierls in 1995. Peierls was Heisenberg’s student, perhaps his most illustrious one, well known to be the father of the atomic bomb, see Frank Close FrankClose1 . In writing this commentary on Peierls’s paper of Physics World in 1991 Peierls1991 I thought it might be useful to recall some memories and discussions I had with Rudy, as we called him in Oxford, pertaining in particular to fundamental questions in quantum mechanics, as a tribute to both men and as a prelude to my views on his response to Bell Bell1990 . Heisenberg was of course famous for his discovery of matrix mechanics in 1925 and as a student Peierls had learnt a lot from him and his way of doing physics and also from Pauli who had many anecdotes which Peierls recalled to me on occasions over the dinner table. Concerning quantum mechanics, I was always interested in discussions with him on the Landau and Peierls (1930) Peierls-Landau1930 paper on the uncertainty in the measurement of fundamental quantities imposed by relativity (finite speed of light $c$), which is the Landau-Peierls expression for the relativistic Heisenberg uncertainty principle: $\Delta p\Delta t\sim\hbar/c$. This would imply amongst other things that a precise measurement for the momentum say of the electromagnetic field i.e. the photon (which by default is a relativistic object) would require an infinite amount of time. This result came as such a shock to Niels Bohr that both young men, Peierls and Landau, were detained in Copenhagen in 1930, unable to leave until Bohr was satisfied that a resolution has been found. In fact, it was not until 1950 that Bohr and Rosenfeld finally published their paper Bohr-Rosenfeld 1950 , which was a masterpiece on measurement analysis, which Bohr insisted must be conducted with classical apparatus. The psychological effects on both Peierls and Landau must have been deep and also on Heisenberg who later became involved Bohr-Rosenfeld 1950 . Heisenberg later changed his own earlier views on quantum mechanics and adopted the S matrix theory for particle interaction processes Heisenberg1938 , as the only description that has observable physical meaning. As for Peierls, my feeling was that he was completely broken down by Bohr (a very persistent and persuasive man) while Landau may not have been so. Landau never accepted quantum field theory as a logically consistent theory due to infinities but partly due to the Landau-Peierls results of 1930 Peierls-Landau1930 ; Landau-Lifshitz4 . Prior to this they had also together penned their paper on quantum electrodynamics in configuration space, proposing for the first time a Schrödinger equation for light quanta Peierls-Landau21930 . The paper inspired Enrico Fermi to use their methodology for the theory of beta decay.